OpenStudy (anonymous):

Solve the equation: \[\frac{ x }{ 3}+\frac{ x }{ 4 }=1\]. Round the solution to the nearest tenth. A. 1.7 B. 2.0 C. 3.2 D. 1.2

OpenStudy (anonymous):

you have a choice, you can add and get \[\frac{4x+3x}{12}=1\] them solve \[\frac{7x}{12}=1\] or you can multiply both sides by \(12\) and solce \[3x+4x=12\] you get the same answer either way

OpenStudy (anonymous):

the answer is A ?

OpenStudy (anonymous):

it is \[x=\frac{12}{7}\] whatever that is

OpenStudy (anonymous):

A :)

OpenStudy (anonymous):

yeah that is closest

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw